# On the Use of Large Language Models in Mutation Testing

### v2

@@ Abstract.
Mutation testing is a foundation approach in the software testing field, based on automatically seeded small syntactic changes, known as mutations. The question of how to generate high-utility mutations, to be used for testing purposes, forms a key challenge in mutation testing literature. Large Language Models (LLMs) have shown great potential in code-related tasks but their utility in mutation testing remains unexplored. To this end, we systematically investigate the performance of LLMs in generating effective mutations w.r.t. to their usability, fault detection potential, and relationship with real bugs. In particular, we perform a large-scale empirical study involving six LLMs, including both state-of-the-art open- and closed-source models, and 851 real bugs on two Java benchmarks (i.e., 605 bugs from 12 projects of Defects4J 2.0 and 246 bugs of ConDefects). We find that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs, which leads to approximately 19% higher fault detection than current approaches (i.e., 93% vs. 74%). Nevertheless, the mutants generated by LLMs have worse compilability rate, useless mutation rate, and equivalent mutation rate than those generated by rule-based approaches. This paper also examines alternative prompt engineering strategies and identifies the root causes of uncompilable mutations, providing insights for researchers to further enhance the performance of LLMs in mutation testing.
Mutation Testing, Mutation Generation, Large Language Models
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Automated Software Engineering; 27 October - 1 November, 2024; Sacramento, CA, USAisbn: 978-1-4503-XXXX-X/XX/XXjournal: JACMjournalvolume: 1journalnumber: FSEarticle: 1publicationmonth: 6ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Mutation testing is one of the most widely-studied software testing techniques (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019). It is performed by producing program variants by making simple syntactic changes (i.e., mutations) in the source code of the origin program (i.e., mutants). These variants form the objectives of testing, in a sense requesting tests to trigger different behaviors on the original and the mutant programs. A mutant is called killed or live depending on whether the test execution leads to a different program output (observable behavior) from that of the original program or not. The proportion of the killed mutants over the entire set of mutants forms a test adequacy metric that is called mutation score.
Since mutations are the objectives of testing, the effectiveness of the method strongly depends on the set of mutants that one is using. Traditional techniques use simple syntactic transformations (rules), called mutation operators, that introduce syntactic changes in every code location that they apply (Jia and Harman, 2010; Papadakis et al., 2019). We call these approaches as rule-based approaches.
Unfortunately, rule-based approaches produce large numbers of redundant mutants, i.e., mutants that do not contribute to the testing process (Papadakis et al., 2016; Ammann et al., 2014) and thus exacerbate the method’s computational overheads as well as the human effort involved in examining live mutants. To address this issue, researchers leverage deep learning approaches to form mutation operators based on the project’s history (Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022; Chen et al., [n. d.]), with the hope that these will produce few but effective mutations. While interesting, these approaches still have limitations in producing effective mutations (Ojdanic et al., 2023b), perhaps due to the fact that they are trained on a limited number of data instances.
To date, the question of how to generate powerful mutants remains open due to the following challenges: (1) Forming effective mutations. To be effective, mutations need to be syntactically correct, which is not the case of most of the existing learning-based methods, and be likely fault revealing (Chekam et al., 2020; Wen et al., 2019), i.e., have a good potential to couple with real faults. While existing methods are effective, there is still significant room for improvement (Degiovanni and Papadakis, 2022). (2) Forming natural mutations. Mutations need to align with natural coding patterns and practices so that they align with developer expectations, i.e., making them feel that they are worth being addressed (Beller et al., 2021). It is a challenge to ensure that the mutated code is not only syntactically correct but also aligns well with the realistic, common coding patterns and practices observed among developers. (3) Forming diverse and killable mutations. Determining whether a mutation is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021). At the same time, mutations need to be diverse and cover the entire spectrum of code and program behaviors.
This paper explores whether Large-Language Models (LLMs), trained on big code and can produce human-like code (Chen et al., 2021), could deal with these challenges. Following existing studies (Guo et al., 2024; Ma et al., 2024; Schäfer et al., 2023), we designed a default prompt that contains task instructions, the code element to be mutated, the corresponding method code, and a few-shot examples sampled from real-world bugs. We carried out a large-scale evaluation involving six LLMs (three closed-source commercial and three open-source ones), 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014a) and 246 bugs from ConDefects (Wu et al., 2024)), and compared with existing mutation testing techniques and tools.
@@
Our evaluation includes both quantitative and qualitative analysis of the mutations generated by LLMs, comparing them with existing approaches in terms of cost, usability (i.e., compilability rate, useless mutation rate, and equivalent mutation rate, AST node diversity), and behavior similarity with real bugs (i.e., real bug detection rate, coupling rate, and Ochiai Coefficient). Additionally, we study how different prompt engineering strategies and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), StarChat-$\beta$-16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs. Overall, our experiments generate about 7.8 million mutations and execute the origin test suites 5.7 million times.
@@
Our results reveal that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs. This leads to higher fault detection rate than traditional approaches. For example, the mutants generated by GPT-4o lead to a fault detection rate of 93.4%, while the mutants generated by LEAM, PIT, and Major are 71.7%, 51.3%, and 74.4%, respectively. Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, useless mutation rate, and equivalent mutation rate. In particular, Major excels with a 98.33% compilability rate, 0% useless mutations, and 0.59% equivalent mutants, while GPT4o exhibits a compilability rate of 75.6%, a useless mutation rate of 7.8%, and an equivalent mutation rate of 1.2%, respectively.
@@
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that GPT-4o-Mini and DeepSeek exhibit the greatest diversity, newly introducing 57 different AST node types. In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 31. The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutations.
@@
We also analyzed the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall in 9 compilation error types with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, which could enhance the effectiveness of mutation testing using AI models.
We wrapped up our LLM-based method into a mutation tool called Wukong. Based on the experimental data, we have also built a dataset that contains high-quality LLM-generated mutants that can be used not only in mutation testing but also in other bug-seeding applications, such as fault localization and fault prediction. The implementation of Wukong and the experimental data are available for open science (our, 2024).
@@
  • We investigate the applicability of LLMs in mutation testing. We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our findings indicate that GPT-4 models excel in generating mutations that closely mimic real bugs.
  • We compare different prompts, and find that few-shot learning with suitable code context archives best performance.
  • We analyze the error types of non-compilable mutations and figure out that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
@@
2. Background and Related Work
Mutation Testing and Mutation Generation.  Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978) one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017). Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014b; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to rule-based. For example, one popular mutation operator is to change one arithmetic binary operator with another one (e.g., $a+b\mapsto a-b$). The majority of existing mutation testing approaches are rule-based, such as PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanch (Schuler and Zeller, 2009), Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), AccMut (Wang et al., 2017), WinMut (Wang et al., 2021), etc. Then researchers proposed using deep learning techniques to generate mutations. For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion. However, these approaches suffer from the challenges in generating syntactic correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations. Perhaps the closest to our work is that of Degiovanni et al. who proposed $\mu$Bert (Degiovanni and Papadakis, 2022). $\mu$Bert generates mutation by masking the code elements to be mutated and then predicting the mutation via Bert (Devlin et al., 2018).
Several recent studies have also explored the use of LLMs for mutation generation. Tip et al. concurrently worked on using LLMs for mutation generation for JavaScript, but their approach neither included a comprehensive comparison with existing techniques nor did they utilize only three open-source LLMs (Tip et al., 2024). Similarly, Endres et al. employed large models for mutation generation to aid deduction (Endres et al., 2024), while Deb et al. used GPT as one of the baselines in their approach (DEB et al., 2024). In contrast to their projects, our work systematically explores a variety of LLMs and prompt designs, offering a thorough comparison with state-of-the-art approaches.
LLMs for Code.  Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023). Existing LLMs are mostly based on Transformer (Vaswani et al., 2017) and are trained by a large corpus of data, including texts and source codes. LLMs are driven by instructions in natural languages, i.e., prompts, which guide the models to perform target tasks and are crucial to performance. The design of prompts plays a crucial role in obtaining accurate and useful responses from these models.
@@ 3.1. Overview and Research Questions
Our paper intends to thoroughly investigate the capability of existing LLMs in mutation generation. We design research questions covering the evaluation of the default setting of our study, the impact of prompt engineering strategies and models, and root cause analysis for underperforming aspects. Specifically, we aim to answer the following five research questions.
  • RQ1: How do the LLMs perform in generating mutations in terms of cost and usability?
  • RQ2: How do the LLMs perform in generating mutations in terms of behavior similarity with real bugs?
  • RQ3: How do different LLMs affect the performance in generating mutations?
  • RQ4: How do different prompts affect the performance in generating mutations?
  • RQ5: What are the root causes and compilation error types for the generation of non-compilable mutations?

### v3

# A Comprehensive Study on Large Language Models for Mutation Testing

@@ Abstract.
Large Language Models (LLMs) have recently been used to generate mutants in both research work and in industrial practice. However, there has been no comprehensive empirical study of their performance for this increasingly important LLM-based Software Engineering application. To address this, we report the results of a comprehensive empirical study over six different LLMs, including both state-of-the-art open- and closed-source models, on 851 real bugs drawn from two different Java real-world bug benchmarks. Our results reveal that, compared to existing rule-based approaches, LLMs generate more diverse mutants, that are behaviorally closer to real bugs and, most importantly, with $90.1\%$ higher fault detection. That is, 79.1% (for LLMs) vs. 41.6% (for rule-based); an increase of $37.5$ percentage points. Nevertheless, our results also reveal that these impressive results for improved effectiveness come at a cost: the LLM-generated mutants have worse non-compilability, duplication, and equivalent mutant rates by $36.1$, $13.1$, and $4.2$ percentage points respectively. These findings are immediately actionable for both research and practice. They allow practitioners to have greater confidence in deploying LLM-based mutation, while researchers now have a baseline for the state-of-the-art, with which they can research techniques to further improve effectiveness and reduce cost.
Empirical Study, Mutation Testing, Mutation Analysis, Mutation Generation, Large Language Models, Coupling
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/XX/XXjournal: TOSEMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
For many years, mutation testing has been the topic of scientific research (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019). Nevertheless, until the last five years, it had made little penetration into the industrial sector. In those most recent five years, it has become increasingly widely deployed in industry. For example, large tech sector companies such as Meta (Foster et al., 2025; Beller et al., 2021) and Google (Petrović et al., 2021) reported on results from deployment of mutation testing at their companies. In particular, results recently reported by Meta (Foster et al., 2025) highlighted the pivotal role played by Large Language Models (LLMs) in the successful deployment of mutation testing at scale in industry.
This recent trend towards LLM-based mutant generation traces back to work on earlier learning-based approaches, which leverage deep learning models to generate mutations by learning patterns from historical bug fixes (Beller et al., 2021; Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022; Chen et al., 2025). These methods aim to introduce mutations that better resemble real faults, potentially increasing their relevance for testing.
However, despite increasing interest in LLM-based mutation testing in both research and practitioner communities, there is to date, no large-scale comprehensive empirical study. Without such a study, we have no basis on which to deploy an industrial scale with confidence. Furthermore, the research community currently has no reliable baseline results for the state-of-the-art effectiveness and efficiency of LLM-based mutation testing compared to more traditional rule-based mutation generation techniques (Jia and Harman, 2010; Papadakis et al., 2019).
The generation of mutants, using any technique is challenging because of the following three important issues:
  1. (1)
    Generating effective mutations: The ultimate goal of mutation testing is to lead to strong tests that uncover real faults, making it critical that generated mutants closely couple with real faults (Chekam et al., 2020; Wen et al., 2019). While existing methods have shown promise, there is still significant room for improvement in enhancing their fault-revealing potential (Degiovanni and Papadakis, 2022).
  2. (2)
    Generating valid mutations.: Mutants must be executable against test cases, which allows tests to assess the program’s behavior in the presence of mutations, to judge whether the mutants are killed. Ensuring executable mutations is challenging because mutations must remain syntactically and semantically valid, respect type constraints, preserve control and data flow, avoid breaking external dependencies, prevent unintended side effects, and align with test case assumptions to ensure meaningful evaluation (Jia and Harman, 2010).
  3. (3)
    Generating diverse and killable mutations: Determining whether a mutation is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021). At the same time, mutations need to be diverse and cover the entire spectrum of code and program behaviors.
This paper reports on how well Large-Language Models (LLMs) handle these challenges, compared to more traditional rule-based approaches. We conducted a comprehensive comparison between six LLMs (three closed-source commercial and three open-source ones) and four existing mutation testing techniques with 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014a) and 246 bugs from ConDefects (Wu et al., 2024)). Overall, our experiments required the generation of 611,100 mutations.
@@
Our evaluation includes both quantitative and qualitative analysis of the mutations generated by LLMs, comparing them with existing approaches in terms of behavior similarity with real bugs (i.e., real bug detection rate, average Ochiai coefficient score, the number of high-Ochiai score bugs, coupling rate, and coupling categorization) as well as mutant validity (i.e., compilability rate, duplicate mutation rate, equivalent mutation rate, and efficiency). Additionally, we study how different prompt engineering strategies and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), StarChat-$\beta$-16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs.
@@
Our results reveal that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs. This leads to a higher fault detection rate than traditional approaches. For example, the mutants generated by GPT-4o-Mini lead to a fault detection rate of 90.8%, while the mutants generated by LEAM, PIT, and Major are 63.9%, 40.1%, and 66.9%, respectively. Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, duplicate mutation rate, and equivalent mutation rate. In particular, Major excels with a 97.6% compilability rate, 0% duplicate mutations, and 2.0% equivalent mutants, while GPT-4o exhibits a compilability rate of 76.4%, a duplicate mutation rate of 7.8%, and an equivalent mutation rate of 3.1%, respectively.
@@
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that GPT-4o-Mini and DeepSeek exhibit the greatest diversity, newly introducing 49 different AST node types. In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 21. The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutations. Moreover, the distributions of AST edit distance of mutations compared to the original code also present that LLMs generate more balanced and expressive mutations. For instance, almost 70% of the mutations of Major have the AST edit distance of one, while for GPT-4o, only 50% of its mutations have the same distance.
@@
We also categorise the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall into 9 compilation error types with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. Moreover, we further analyze the code context of the non-compilable mutations generated by LLM-based approaches, and find that Method Reference and Method Invocation are the most frequent code features. This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, which could enhance the effectiveness of mutation testing using AI models.
Our implementation and the experimental data are available at our repo: https://github.com/da-123-snake/Mutantion-Generation-Study.
@@
  • We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our findings indicate that LLMs excel in generating diverse mutations that closely mimic real bugs.
  • We analyze the error types of non-compilable mutations and determine that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
@@
The remainder of this paper is organized as follows. Section 2 presents the design of our study, including the research questions, studied approaches, datasets, and metrics. Section 3 presents the empirical evaluation of these generation approaches and answers to each research question. Section 4 discusses the influences of different experimental setups and the implications summarized from the findings. Section 5 presents the threats to the validity of our study. Section 6 discusses the related work. Section 7 concludes this work.
@@ 2. Study Design
In this section, we introduce how our study is designed and conducted.
@@ 2.1. Overview and Research Questions
This paper intends to evaluate existing mutation generation approaches by answering the following research questions.
  • RQ1 (Effectiveness) How well do mutations generated by LLMs mimic real bugs compared to existing methods? This research question investigates the effectiveness of LLM-generated mutations in representing real bugs. Specifically, it evaluates whether these mutations exhibit similar characteristics to real faults in terms of detection capability, behavioral similarity, and diversity. This RQ is divided into the following sub-RQs:
    • RQ1.1: (Detectability) To what extent do bug-triggering tests successfully detect LLM-generated mutations?
    • RQ1.2 (Coupling) To what extent do LLM-generated mutations couple with real bugs?
    • RQ1.3 (Diversity) How diverse are LLM-generated mutations compared to traditional approaches?
  • RQ2 (Validity) How do LLMs compare to existing approaches in terms of validity? This research question explores the validity of mutations generated by LLMs compared to traditional approaches. Validity is assessed based on key factors such as compilability, duplication, and equivalence of generated mutations. By analyzing these aspects, we aim to determine whether LLMs produce valid and meaningful mutations that contribute to practical mutation testing. This RQ is divided into the following sub-RQs:
    • RQ2.1 (Compilability) What percentage of generated mutations are compilable across different approaches?
    • RQ2.2 (Duplication) What percentage of the generated mutations are duplicates of the original code across different approaches?
    • RQ2.3 (Equivalence) What percentage of generated mutations are equivalent across different approaches?
  • RQ3 (Efficiency) What are the time costs of different mutation generation approaches? This research question investigates the time costs of different mutation generation approaches. Time costs are associated with the efficiency and practicality of mutation generation approaches.
  • RQ4 (Compilation Error) What are the types of compilation errors of the non-compilable mutations of different approaches? This RQ aims to categorize and analyze the types of compilation errors that arise from non-compilable mutations generated by different approaches. Understanding these errors can guide mutation generation approaches to enhance their practical usability.
2.2. Studied Mutation Generation Approaches
Our study is conducted in Java because: 1) it is one of the most widely used programming languages, shown as the TIOBE ranking [^1] https://www.tiobe.com/tiobe-index/; 2) the majority of existing mutation generation approaches is for Java. We categorize these approaches into two groups: traditional approaches (i.e., those developed before the emergence of LLMs) and LLM-based approaches (i.e., approaches that leverage LLMs and are designed for interaction in human-like conversations), as shown below. Table 1 summarizes the mutation generation approaches covered by this study.
Table 1. Studied Mutation Generation Approaches
Type Name Description
Traditional PIT Widely studied mutation testing framework with 29 mutation operators.
Major Widely studied mutation testing framework with 11 mutation operators.
LEAM Custom-trained models learning from a large corpus of real bugs and patches.
$\mu$Bert Mutation generation with the pre-trained model CodeBERT.
LLM-based GPT-3.5 GPT-3.5-Turbo, whose based model is GPT, trained on data before 2021/09, released in 2021/03.
GPT-4o The based model is GPT, trained on data before 2023/10, released in 2024/05.
GPT-4o-M GPT-4o-Mini, whose based model is GPT, trained on data before 2023/10, released in 2024/07.
SC-16b StarChat-$\beta$-16b, whose based model is StarCoder, released in 2023/06.
CL-13b CodeLlama-Instruct-13b, whose based model is Llama, released in 2023/08.
DS-236b DeepSeek-Coder-V2-236b, whose based model is DeepSeek, released in 2024/07.
2.2.1. Traditional Approaches
For Java, there exist several mutation testing frameworks, such as JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), PIT (Coles et al., 2016), and Major (Just et al., 2011). As some tools are no longer maintained, following existing studies (Kaufman et al., 2022; Ojdanic et al., 2023; Kushigian et al., 2024), we employ PIT (Coles et al., 2016), Major (Just et al., 2011), LEAM (Tian et al., 2022), and $\mu$Bert (Degiovanni and Papadakis, 2022). Major and PIT are widely studied rule-based approaches, while LEAM and $\mu$Bert involve deep learning models.
PIT: PIT (Coles et al., 2016) is a widely used mutation testing tool on the Java bytecode level, which supports 29 well-designed mutation operators and four groups of active mutation operator sets. In our study, we employ PIT 1.14.4 and leverage the ALL group which activates all operators.
Major: Major (Just et al., 2011) is another high-impact mutation testing framework, that supports nine operators. In our study, we use Major 1.3.5 and involve all mutation operators.
LEAM: LEAM (Tian et al., 2022) is the state-of-the-art custom-trained-model-based approach, which employs sequence-to-sequence models to generate mutations. Its model is trained on a large corpus of real-world bugs and their corresponding patches, allowing it to learn patterns for generating more effective mutations.
$\mu$Bert: $\mu$Bert (Degiovanni and Papadakis, 2022) is the state-of-the-art mutation generation approaches based on the small-scale pre-trained model, Bert (Devlin et al., 2018). Equipped with the pre-trained model, $\mu$Bert transforms the task of mutation generation to code completion under the surrounding context.
2.2.2. LLM-based Approaches
Our study involve multiple kinds of LLMs, including both open-source and closed-source models.
LLMs: Our study also comprehensively covers the most widely studied LLMs. Specifically, we select a representative set of LLMs with code capabilities both from commercial closed-source and open-source models. Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from the GPT family, including GPT-3.5 (GPT-3.5-Turbo) (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-M (GPT-4o-Mini) (Achiam et al., 2023)), and open-source LLMs, including SC-16b (StarChat-$\beta$-16b), CL-13b (CodeLlama-Instruct-13b), and DS-236b (DeepSeek-Coder-V2-236b). The details of these models are shown as Table 1. Note that OpenAI has not officially released the parameter size of their closed-source models. The parameter size of the open-sourced models we studied ranges from 13 billion to 236 billion.